OpenStudy (anonymous):

Linda had been collecting stamps for 5 weeks and had 25 stamps in her stamp book. Her book holds 64 stamps. She then began purchasing 3 stamps each week. What was the total number of weeks that it took Linda to fill her stamp book? _____weeks

OpenStudy (alexandervonhumboldt2):

64-25=39 number left 39/3=13 weeks

OpenStudy (alexandervonhumboldt2):

13+5=18 weeks she coolect them

OpenStudy (anonymous):

thanks